Pepperoni:

When Jeremy surveyed thirty 7th-grade students at the community pool, he found that 2/15 had not read at least one book during July. He knows there are about 200 7th-grade students in his community, so he infers that about 40 of them did not read any books in July. Is that a good inference? Explain.

whybruh:

No. 40/200 = 1/5 2/15 = 1/7.5 So I would say that it is not that good of a reference. You may find a way to explain it yourself though.

Pepperoni:

ok

whybruh:

ok